# A fast Bellman algorithm for the real Monge–Ampère equation

### v2

@@ 2. Background
for any Borel subset $E\subset\Omega$, where $\operatorname{Leb}$ denotes the Lebesgue measure, $\partial u(E)=\cup_{x\in E}\partial u(x)$, and $\partial u(x)$ is the set of all slopes of the supporting hyperplanes to the graph of $u$ at $x$. If $u\in C^{2}$ and $D^{2}u$ is positive semi-definite, then $\mu_{u}(E)=\det(D^{2}u)$. We should mention, that in contrast to the real setting, it is impossible to define the complex Monge–Ampère operator for every plurisubharmonic function. We refer to the monograph by Guedj and Zeriahi [17] for the theory of complex Monge–Ampère equations.
@@
Furthermore, if $\Omega$ is strictly convex (or uniformly convex in the unbounded setting) with $C^{k+2,\alpha}$ boundary, $0<\phi\in C^{k+2,\alpha}(\partial\Omega)$ and $f\in C^{k,\alpha}(\bar{\Omega})$ for some integer $k\geq 2$ and some $10<\alpha<$ 1, then the Dirichlet problem (2.1) has a unique convex solution $u\in C^{k+2,\alpha}(\bar{\Omega})$, see [9, 21, 22].
@@
There have been a number of suggested numerical methods for solving the Dirichlet problem for the Monge–Ampère operator. As far as we are aware, the first approach was due to Oliker and Prussner (1988) [27] who proposed a direct discretization of the Aleksandrov interpretation of the Monge–Ampère operator. Since then, a number of other approaches to solving Monge–Ampère equations have been tried: finite difference discretization, often using wide stencils to obtain stable and consistent methods [5, 6, 11, 12, 26], variational methods [4, 19], fixpoint methods [5, 18] as well as finite elements [20, 23] only to mention a few.
@@ 3. Bellman’s principle
Our method relies on the following linear algebra result, sometimes called Bellman’s principle. In the setting of complex Monge–Ampère equations, this idea was introduced by Gaveau [15] and has later found several theoretical applications. We will formulate the result in the real setting and for convenience, we provide a proof of Bellman’s principle, similar to the approach taken by Wiklund [29].
@@
To this end we observe that for any symmetric and positive definite matrix $B$, we have
@@
for which there are highly efficient numerical algorithms. Let us mention that with this approach we have the following monotonicity result.
Proposition 3.3.
Let $\Omega\subset\mathbb{R}^{n}$ be a convex domain with $C^{2,\alpha}$ boundary, $\phi\in C^{2,\alpha}(\partial\Omega)$, $f^{1/n}\in C^{0,\alpha}(\Omega)$ and let $B\in C^{0,\alpha}(\Omega)$ be a $\mathcal{B}$-valued function on $\Omega$. Assume that the solution to the above Dirichlet problem (Eq. (3.2)) is convex. Then $u\geq u_{\text{MA}}$ where $u_{\text{MA}}$ denotes the solution to the corresponding Monge–Ampère equation (2.1).
Proof.
By standard regularity theory for elliptic PDEs, $u\in C^{2,\alpha}$, so by Corollary 3.2, $\operatorname{MA}(u)\leq f$. Hence, by the comparison principle for the real Monge–Ampère operator (see for example [24, Theorem 3.21]), it follows that $u\geq u_{\text{MA}}$. ∎
Unfortunately, it seems like there are no known effective characterizations of when the solution to Equation (3.2) is convex, even in the case when $B\equiv I$, i.e. for Poisson’s equation. Our numerical experiments suggest that our proposed algorithm (as described in Section 5) produces a decreasing sequence of approximants converging to the solution of (2.1), at least once one of the approximants is convex. So far, we have not been able to prove this in full generality, but under suitable technical conditions and the assumption that the method produces a sequence of strictly convex approximants, we can show convergence, see Theorem 4.1.
@@ 4. Proof of convergence
Let $\alpha\in(0,1)$ and let $\Omega\subset\mathbb{R}^{2}$ be a strictly convex, bounded, $C^{2,\alpha}$ domain, $\phi\in C^{2,\alpha}(\bar{\Omega})$ and $f\in C^{0,\alpha}(\bar{\Omega})$ where $f>0$ on $\bar{\Omega}$. Assume that there exist strongly convex solutions $\{u_{k}\}_{k\geq 0}\subset C^{2,\alpha}(\bar{\Omega})$ to the following linear Dirichlet problems
(4.1) $\begin{cases}\operatorname{tr}(B_{k}(x)D^{2}u_{k}(x))=2\sqrt{f(x)},&x\in\Omega,\\ u_{k}(x)=\phi(x),&x\in\partial\Omega,\end{cases}$
where $B_{k}=\sqrt{\det D^{2}u_{k}}\,(D^{2}u_{k})^{-1}$ for $k\geq 1$ and $B_{0}$ is a $\mathcal{B}$-valued function with coefficients $(B_{0})_{ij}\in C^{0,\alpha}(\bar{\Omega})$. Moreover, denote by $\lambda_{k},\Lambda_{k}$ the eigenvalues of $D^{2}u_{k}$ and assume that there exists $\epsilon>0$ such that $\epsilon\leq\lambda_{k},\Lambda_{k}$ for all $k$. Then, there exists a subsequence of the solutions $\{u_{k}\}_{k\geq 0}$ converging uniformly to a convex function $\tilde{u}\in C^{1,1}(\bar{\Omega})$ solving the Monge–Ampère equation (2.1).
@@
Now, let us investigate solvability and well-definiteness of the next Dirichlet problem in the sequence. We would like to generalise this argument for any $k\geq 0$. To this end, assume the existence of the solution $u_{k}\in C^{2,\alpha}(\bar{\Omega})$.
Observe that $B_{k+1}\coloneqq\sqrt{\det D^{2}u_{k}}(D^{2}u_{k})^{-1}$ is not necessarily well-defined, nor is the next differential operator $L_{k+1}(v)\coloneqq\operatorname{tr}(B_{k+1}D^{2}v)$ elliptic. Consequently, the solution to the corresponding Dirichlet problem is not guaranteed.
However, strong convexity of $u_{k}$ renders the next differential operator $L_{k+1}$ well-defined and elliptic. To this end, note that $D^{2}u_{k}\in C^{0,\alpha}(\bar{\Omega})$ and thus $\det(D^{2}u_{k})\in C^{0,\alpha}(\bar{\Omega})$. Hence, both eigenvalues are bounded from above. Now, if $\epsilon\leq\lambda_{k},\Lambda_{k}$ it follows that $D^{2}u_{k}$ is positive definite and hence invertible with a positive determinant. Thus $B_{k+1}$ is well-defined and by construction it is also positive definite. Furthermore, the regularity of the solution, i.e. $u_{k}\in C^{2,\alpha}(\bar{\Omega})$ results in $(B_{k})_{ij}\in C^{0,\alpha}(\bar{\Omega})$. Thus $L_{k+1}$ is elliptic with the necessary regularity of coefficients for the corresponding Dirichlet problem to have a unique solution $u_{k+1}\in C^{2,\alpha}(\bar{\Omega})$ (by another application of Theorem 6.14 in Gilbarg and Trudinger).
@@
We begin with defining $\nu_{k}\coloneqq\det D^{2}u_{k}(x)$ and noting that $\{u_{k}\}_{k\geq 0}$ are solutions to the following non-linear Dirichlet problems:
@@
Moreover, the existence of a uniform lower bound of the eigenvalues $\lambda_{k},\Lambda_{k}$ implies the existence of a uniform upper bound.
Indeed, recall that since $B_{k+1}=\sqrt{\det D^{2}u_{k}}\,(D^{2}u_{k})^{-1}$, it is positive definite with $\det B_{k+1}=1$, for every $k\geq 0$, and we can use the Corollary to Bellman’s principle (Corollary 3.2) to conclude that
(4.3) $2\epsilon\leq 2\sqrt{\det D^{2}u_{k}}=\operatorname{tr}{B_{k+1}D^{2}u_{k}}\leq\operatorname{tr}{B_{k}D^{2}u_{k}}=2\sqrt{f}\text{ on }\bar{\Omega}.$
Thus,
(4.4) $\epsilon\leq\lambda_{k},\Lambda_{k}\leq\|\sqrt{f}\|_{C(\bar{\Omega})}/\epsilon,\qquad\text{for $k\geq 0$.}$
Consequently, $\{\nu_{k}\}_{k}$ can be interpreted as finite Borel measures with $\nu_{k}(\Omega)\leq(\|f\|_{C(\Omega)}/\epsilon)^{2}$ and $\{u_{k}\}_{k}$ as Aleksandrov solutions to (4.2). Moreover, there exists a finite Borel measure $\nu$ and a subsequence $\{\nu_{k}\}_{k\geq 0}$ (after re-enumeration) such that $\nu=\mathop{w^{*}\text{-}\lim}_{k\to\infty}\nu_{k}$. By compactness of Aleksandrov solutions to the Dirichlet problem (4.2) ([24, Theorem 3.35]), we can conclude that there exists $\tilde{u}\in C(\bar{\Omega})$ and a subsequence $\{u_{k}\}_{k\geq 0}$ (after re-enumeration) such that $u_{k}\to\tilde{u}$ uniformly and $\tilde{u}$ satisfies the following
@@
It remains to show that $\tilde{u}$ solves the initial Monge–Ampère equation (2.1) by showing that $f=\nu$ as measures. To this end, we observe that all partial derivatives $\{(D^{2}u_{k})_{ij}\}_{k\geq 0}$ are uniformly bounded from above and below.
Indeed, we can use Rayleigh quotients to conclude that $\{(D^{2}u_{k})_{ii}\}_{k\geq 0}$ are bounded by the eigenvalues. The existence of uniform lower and upper bounds follows from Equation (4.4). Now, using $\det D^{2}u_{k}=(D^{2}u_{k})_{ii}(D^{2}u_{k})_{jj}-(D^{2}u_{k})^{2}_{ij}$ and Equation (4.3), we see that $\{(D^{2}u_{k})_{ij}\}_{k\geq 0}$ are also uniformly bounded. Thus we can choose a subsequence $\{u_{k}\}_{k\geq 0}$ (after re-enumeration) such that $u_{k}\to\tilde{u}$ uniformly and $\{(D^{2}u_{k})_{ii}\}_{k\geq 0}$, $\{(D^{2}u_{k})_{jj}\}_{k\geq 0}$ and $\{(D^{2}u_{k})_{ij}\}_{k\geq 0}$ have a weak limit.
To show that $f=\nu$ in measures, we recall Equation (4.1) and $\operatorname{tr}(B_{k+1}D^{2}u_{k})=2\sqrt{\det(D^{2}u_{k})}\coloneqq 2\sqrt{\nu_{k}}$ and establish the following
$2\sqrt{f}$ $=\operatorname{tr}(B_{k}D^{2}u_{k})=\operatorname{tr}(B_{k}-B_{k+1})D^{2}u_{k}+\operatorname{tr}(B_{k+1}D^{2}u_{k})$
$=\underbrace{\operatorname{tr}(B_{k}-B_{k+1})D^{2}u_{k}}_{(*)}+\underbrace{2\sqrt{\nu_{k}}}_{(**)}.$
Furthermore, recall that $B_{k}=\sqrt{\det D^{2}u_{k-1}}(D^{2}u_{k-1})^{-1}$, so Cramer’s rule gives
$\begin{split}(B_{k})_{11}&=(D^{2}u_{k-1})_{22}/\sqrt{\det D^{2}u_{k-1}},\\ (B_{k})_{22}&=(D^{2}u_{k-1})_{11}/\sqrt{\det D^{2}u_{k-1}},\\ (B_{k})_{12}&=-(D^{2}u_{k-1})_{21}/\sqrt{\det D^{2}u_{k-1}}.\end{split}$
Consequently, $(*)\xrightarrow{w^{*}}0$ as $k\to\infty$ since the denominators are uniformly bounded away from 00 by (4.3). Furthermore ()(**)( ∗ ∗ ) is also uniformly bounded by Equation (4.3). This yields
$f-\nu=\mathop{w^{*}\text{-}\lim}_{k\to\infty}(f-\nu_{k})=\mathop{w^{*}\text{-}\lim}_{k\to\infty}(\sqrt{f}-\sqrt{\nu_{k}})(\sqrt{f}+\sqrt{\nu_{k}})=0.$
Since $u_{k}$ converge uniformly and all the second-order derivatives are uniformly bounded, it follows from the Arzelà–Ascoli theorem that $\tilde{u}\in C^{1,1}(\bar{\Omega})$. ∎

### v3

@@ 2. Background
for any Borel subset $E\subset\Omega$, where $\operatorname{Leb}$ denotes the Lebesgue measure, $\partial u(E)=\cup_{x\in E}\partial u(x)$, and $\partial u(x)$ is the subgradient of $u$, i.e. the set of all slopes of the supporting hyperplanes to the graph of $u$ at $x$. If $u\in C^{2}$ and $D^{2}u$ is positive semi-definite, then $\mu_{u}(E)=\det(D^{2}u)$. We should mention, that in contrast to the real setting, it is impossible to define the complex Monge–Ampère operator for every plurisubharmonic function. We refer to the monograph by Guedj and Zeriahi [17] for the theory of complex Monge–Ampère equations.
@@
Furthermore, if $\Omega$ is strictly convex (or uniformly convex in the unbounded setting) with $C^{k+2,\alpha}$ boundary, $0<\phi\in C^{k+2,\alpha}(\partial\Omega)$ and $f\in C^{k,\alpha}(\bar{\Omega})$ for some integer $k\geq 2$ and some $0<\alpha<1$, then the Dirichlet problem (2.1) has a unique convex solution $u\in C^{k+2,\alpha}(\bar{\Omega})$, see [9, 21, 22].
@@
There have been a number of suggested numerical methods for solving the Dirichlet problem for the Monge–Ampère operator. As far as we are aware, the first approach was due to Oliker and Prussner (1988) [28] who proposed a direct discretization of the Aleksandrov interpretation of the Monge–Ampère operator. Since then, a number of other approaches to solving Monge–Ampère equations have been tried: finite difference discretization, often using wide stencils to obtain stable and consistent methods [5, 6, 11, 12, 27], variational methods [4, 19], fixed-point methods [5, 18] as well as finite elements [20, 23] only to mention a few.
After the preparation of an earlier draft of this paper, we learned that a similar method as ours (for periodic boundary values) has been previously proposed by Loeper and Rapetti [26]. Their approach is somewhat different than ours, and derive their method as a Newton type iteration for the non-linear differential operator.
@@ 3. Bellman’s principle
Our method relies on the following linear algebra result, sometimes called Bellman’s principle. In the setting of complex Monge–Ampère equations, this idea was introduced by Gaveau [15] and has later found several theoretical applications, both in the real and the complex setting. We will formulate the result in the real setting and for convenience of the reader, we provide a proof of Bellman’s principle.
@@
To this end, we observe that for any symmetric and positive definite matrix $B$, we have
@@
for which there are highly efficient numerical algorithms.
Unfortunately, it seems like there are no known effective characterizations of when the solution to Equation (3.2) is convex, even in the case when $B\equiv I$, i.e. for Poisson’s equation. Under suitable technical conditions and the assumption that our method produces a sequence of strictly convex approximants, we can show convergence, see Theorem 4.1.
@@ 4. Proof of convergence
Let $\alpha\in(0,1)$ and let $\Omega\subset\mathbb{R}^{n}$ be a strictly convex, bounded, $C^{2,\alpha}$ domain, $\phi\in C^{2,\alpha}(\bar{\Omega})$ and $f\in C^{0,\alpha}(\bar{\Omega})$ with strictly positive values on $\bar{\Omega}$. Assume that there exist solutions $\{u_{k}\}_{k\geq 0}\subset C^{2,\alpha}(\bar{\Omega})$ to the following linear Dirichlet problems
(4.1) $\begin{cases}\operatorname{tr}(B_{k}(x)D^{2}u_{k}(x))=nf^{1/n}(x),&x\in\Omega\\ u_{k}(x)=\phi(x),&x\in\partial\Omega\\ u_{k}(x)\text{ is strongly convex}.\end{cases}$
where $B_{k}=(\det{D^{2}u_{k-1}})^{1/n}(D^{2}u_{k-1})^{-1}=\operatorname{cof}D^{2}u_{k-1}$ for $k\geq 1$ and $B_{0}$ be a $\mathcal{B}$-valued function with coefficients $(B_{0})_{ij}\in C^{0,\alpha}(\bar{\Omega})$. Moreover, denote by $\lambda_{\text{min}}^{k}$ the smallest eigenvalue of $D^{2}u_{k}$ and assume that $\{\lambda_{\min}^{k}\}_{k\geq 0}$ is uniformly bounded away from zero, i.e. there exists $\epsilon>0$ such that $\lambda_{\text{min}}^{k}\geq\epsilon$ for every $k\geq 0$. Then, the sequence of solutions $\{u_{k}\}_{k\geq 0}$ converges uniformly to a convex function $\tilde{u}\in C^{1,1}(\bar{\Omega})$ solving the Monge–Ampère equation (2.1).
@@
Next, let us investigate solvability and well-definiteness of the next Dirichlet problem in the sequence. We would like to generalise this argument for any $k\geq 0$. To this end, assume the existence of the solution $u_{k}\in C^{2,\alpha}(\bar{\Omega})$.
Observe that $B_{k+1}\coloneqq(\det D^{2}u_{k})^{1/n}\,(D^{2}u_{k})^{-1}$ is not necessarily well-defined, nor is the next differential operator $L_{k+1}(v)\coloneqq\operatorname{tr}(B_{k+1}D^{2}v)$ strictly elliptic. Consequently, the solution to the corresponding Dirichlet problem is not guaranteed.
However, strong convexity of $u_{k}$ renders the next differential operator $L_{k+1}$ well-defined and strictly elliptic. Indeed, first note that from $\epsilon\leq\lambda_{\text{min}}^{k}$ it follows that $D^{2}u_{k}$ is positive definite and hence invertible. Recall that $B_{k+1}\coloneqq(\det D^{2}u_{k})^{1/n}(D^{2}u_{k})^{-1}$, by the above $B_{k+1}$ is well-defined and positive definite with $\det B_{k+1}=1$. We can use the Corollary to Bellman’s principle (Corollary 3.2) to conclude that
$n\epsilon\leq n(\det D^{2}u_{k})^{1/n}=\operatorname{tr}{B_{k+1}D^{2}u_{k}}\leq\operatorname{tr}{B_{k}D^{2}u_{k}}=nf^{1/n}\text{ on }\bar{\Omega}.$
Thus,
(4.2) $\epsilon^{n}\leq\det D^{2}u_{k}\leq f\text{ on }\bar{\Omega}.$
Moreover, existence of a uniform lower bound of the eigenvalues $\lambda^{k}_{min}$ implies existence of a uniform upper bound for the largest eigenvalues denoted by $\lambda^{k}_{max}$. Indeed, Eq. (4.2) yields
$\epsilon\leq\lambda^{k}_{min}\leq\lambda^{k}_{max}\leq f/\epsilon^{n-1},\qquad\text{for $k\geq 0$.}$
By diagonalising $B_{k+1}$ we have that the smallest and the largest eigenvalues are uniformly bounded as follows
$\begin{split}\gamma_{min}^{k}&=(\det D^{2}u_{k})^{1/n}/\lambda_{max}^{k}\geq(\epsilon^{n})^{1/n}/(f/\epsilon^{n-1})\geq\epsilon/\|f\|_{C(\bar{\Omega})}\text{ and }\\ \gamma_{max}^{k}&=(\det D^{2}u_{k})^{1/n}/\lambda_{min}^{k}\leq\|f^{1/n}\|_{C(\bar{\Omega})}/\epsilon.\end{split}$
From the lower bound of $\gamma_{min}^{k}$ follows strict ellipticity of $L_{k+1}$. Furthermore, the regularity of the solution, i.e. $u_{k}\in C^{2,\alpha}(\bar{\Omega})$ results in $(B_{k+1})_{ij}\in C^{0,\alpha}(\bar{\Omega})$ as $(B_{k+1})_{ij}$ are obtained by adding, subtracting and multiplying terms $\partial_{ij}^{2}u_{k}$.
Thus $L_{k+1}$ is strictly elliptic with the necessary regularity of coefficients for the corresponding Dirichlet problem to have a unique solution $u_{k+1}\in C^{2,\alpha}(\bar{\Omega})$ (by another application of Theorem 6.14 in Gilbarg and Trudinger).
@@
We begin with showing monotonicity and hence convergence of the sequence of solutions $\{u_{k}\}_{k\geq 0}$.
Let us define $\nu_{k}\coloneqq\det D^{2}u_{k}(x)$ and $\eta_{k}\coloneqq u_{k+1}-u_{k}$. Then
$\operatorname{tr}B_{k+1}D^{2}\eta_{k}$ $=\operatorname{tr}B_{k+1}D^{2}u_{k+1}-\operatorname{tr}B_{k+1}D^{2}u_{k}$
$=nf^{1/n}-\operatorname{tr}(\det D^{2}u_{k})^{1/n}\,(D^{2}u_{k})^{-1}D^{2}u_{k}$
$=nf^{1/n}-n(\det D^{2}u_{k})^{1/n}=nf^{1/n}-n(\nu_{k})^{1/n}.$
Hence $\eta_{k}$ solves the following linear elliptic PDE, the coefficients of which are elements of the matrix $B_{k+1}\coloneqq\operatorname{cof}D^{2}u_{k}$,
$\begin{cases}\operatorname{tr}B_{k+1}D^{2}\eta_{k}=n(f)^{1/n}-n(\nu_{k})^{1/n}\text{ in }\Omega,\\ \eta_{k}=0\text{ on }\partial\Omega.\end{cases}$
Since the operator $L_{k+1}(v)\coloneqq\operatorname{tr}B_{k+1}(D^{2}v)$ is strictly elliptic and Eq. (4.2)) gives $n(f)^{1/n}-n(\nu_{k})^{1/n}\geq 0$, the maximum principle for linear elliptic PDEs ([16, Theorem 3.1]) yields
$u_{k}-u_{k-1}=\eta_{k}\leq\max_{\bar{\Omega}}\eta_{k}=\max_{\partial\Omega}\eta_{k}=0,$
or in other words, for every $k\geq 0$.
(4.3) $u_{k}\leq u_{k-1}\text{ on }\bar{\Omega}.$
It follows from the above and from the comparison principle for the real Monge–Ampère operator (see for example [24, Theorem 3.21]) that $\{u_{k}\}_{k\geq 0}$ is a decreasing sequence bounded from below by $u_{MA}$, the solution to the corresponding Monge–Ampère equation (2.1). Hence, the sequence converges to some function $\tilde{u}\coloneqq\lim_{k\to\infty}u_{k}$. Moreover, for every $k\geq 0$ and $\psi\in C^{\infty}_{0}(\Omega)$
$\int_{\Omega}(\partial_{ij}^{2}u_{k+1})\psi dV=\int_{\Omega}u_{k}(\partial_{ij}^{2}\psi)dV.$
The monotone convergence theorem implies that all second partial derivatives $\{\partial_{ij}^{2}u_{k}\}_{k}$ converge weak-star. Hence, we can define a limit function $\nu\coloneqq\mathop{w^{*}\text{-}\lim}_{k\to\infty}\nu_{k}$. Since $0<\epsilon^{n}\leq\nu,\nu_{k}\leq f$ for all $k\geq 0$ (Eq. (4.2)), we can interpret $\nu_{k}$ and $\nu$ as finite Borel measures. We note that $\{u_{k}\}_{k\geq 0}$ are Aleksandrov solutions to the following non-linear Dirichlet problems:
@@
Next, we use compactness of Aleksandrov solutions to the Dirichlet problem (4.4) ([24, Theorem 3.35]) to conclude that the limit function $\tilde{u}\coloneqq\lim_{k\to\infty}u_{k}\in C(\bar{\Omega})$ and satisfies the following
@@
Since the limit function is continuous and the sequence monotone, Dini’s theorem gives that the convergence is uniform. It remains to show that $\tilde{u}$ solves the initial Monge–Ampère equation (2.1) by showing that $f=\nu$ as measures. To this end, let us note the following
(4.5) $\begin{split}n(f)^{1/n}&=\operatorname{tr}(B_{k+1}D^{2}u_{k+1})=\operatorname{tr}B_{k+1}(D^{2}u_{k+1}-D^{2}u_{k})+\operatorname{tr}B_{k+1}D^{2}u_{k}\\ &=\underbrace{\operatorname{tr}B_{k+1}(D^{2}u_{k+1}-D^{2}u_{k})}_{\coloneqq\tau}+n(\underbrace{\nu_{k}}_{\xrightarrow[]{w^{*}}\nu})^{1/n}.\end{split}$
Hence, it remains to show that $\tau\to 0$ weak-star. Since all second partial derivatives $\partial_{ij}^{2}u_{k}$ converge weak-star, it suffices to uniformly bound terms $(B_{k})_{ij}$. To this end, we first define bilinear forms $(x,y)_{k}\coloneqq(B_{k}x,y)$ for all $x,y\in\mathbb{R}^{n}$. From Rayleigh quotients, we conclude that $\gamma^{k}_{min}\leq\|x\|^{2}_{k}/\|x\|^{2}\leq\gamma^{k}_{max}$ for all $x\in\mathbb{R}^{n}$ and for all $k\geq 0$. Let $\{e_{m}\}_{1\leq m\leq n}$ denote the standard basis in $\mathbb{R}^{n}$ and use the polarization identity to obtain
$(B_{k})_{ij}=(e_{j},e_{i})_{k}=(1/4)(\|e_{j}+e_{i}\|^{2}_{k}-\|e_{j}-e_{i}\|^{2}_{k}).$
Recalling Equation (4), it follows that
(4.6) $\epsilon/\|f\|_{C(\bar{\Omega})}\leq\gamma_{min}^{k}\leq(B_{k})_{ij}\leq\gamma_{max}^{k}\leq\|f^{1/n}\|_{C(\bar{\Omega})}/\epsilon,$
for all $1\leq i,j\leq n$ and $k\geq 0$. Thus, we have that
$\tau\coloneqq\operatorname{tr}B_{k+1}(D^{2}u_{k+1}-D^{2}u_{k})\xrightarrow[]{w^{*}}0.$
By the above and Equation (4.5) we have
$f-\nu=\mathop{w^{*}\text{-}\lim}_{k\to\infty}(f-\nu_{k})=(f^{1/n}-\nu_{k}^{1/n})\sum_{j=0}^{n-1}(-1)^{j}f^{(n-1-j)/n}\nu_{k}^{j/n}=0.$
Thus, $f=\nu$ as measures. Since all second partial derivatives of $\{u_{k}\}_{k\geq 0}$ are bounded, the Arzelà–Ascoli theorem gives that in fact $\tilde{u}\in C^{1,1}(\bar{\Omega})$. ∎